OpenStudy (anonymous):

Please Help! Solve x3 = 125 over 27. 25 over 9 ±25 over 9 5 over 3 ±5 over 3

OpenStudy (anonymous):

what it the cubed root of 27?

OpenStudy (anonymous):

The cubed root of 27 is 3x3x3, right?

OpenStudy (anonymous):

it is true that \(3\times 3\times 3=27\) so the cubed root of 27 is 3 how about the cubed root of \(125\) ?

OpenStudy (anonymous):

5x5x5 = 125, so the cubed root would be 5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what is the cubed root of \(\frac{125}{27}\) ?

OpenStudy (anonymous):

Would it be \[\frac{ 5 }{ 3 }\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you so much!!!